Solve for x. 1/y = 3/5 - 2/x
Answer by jai_kos(139)

Given 1/y = 3/5 - 2/x
Leave the term which contains "x" and bring all the other term to the other side.
We have,
1/y - 3/5 = - 2 /x
MUltiply the above by "x", we get
- 1/ y + 3 / 5 = 2 /x
Simplify the above, we get
- 5 + 3y / 5y = 2 / x
Cross multiply the terms, we get
x ( -5 + 3y ) = 2 ( 5y)
x ( -5 + 3y) = 10 y
x = 10 y / (-5 + 3y)
The above solution will solve for "x"